A harmonic wave is traveling along a rope. It is observed that the oscillator that generates
the wave completes 40.2 vibrations in 33.7 s. Also, a given maximum travels 438 cm alongthe rope in 11.7 s.

What is the wavelength?

Answer in units of cm.
1
Expert's answer
2019-07-30T08:49:59-0400

The frequency of the wave is f = 40.2/33.7 = 1.193 1/s. The speed of the wave - v = 4.38/11.7 = 0.37 m/s The wavelength would be equal to λ = v/f = 0.37 [m/s]/1.193 [1/s] = 0.31 m = 31 cm.